make your own problem .one problem calculating the period and the frequency of spring mass.show your complete solution to the problem.




1
Expert's answer
2021-02-10T18:23:26-0500

"f=\frac{1}{2\pi}\sqrt{\frac{k}{m}}\\f=\frac{1}{2\pi}\sqrt{\frac{100}{1}}=1.59\ Hz\\T=2\pi\sqrt{\frac{1}{100}}=0.628\ s"